# KeyB2: Selecting Key Blocks is Also Important for Long Document Ranking with Large Language Models

### v1

@@ 5.7. Analysis of the Position of Selected Blocks
Across all datasets, over half of the top 8 selected blocks come from the first 5 positions in documents, reflecting a common positional bias where initial content is often more relevant to the query. Additionally, the figures indicate that relevance distributions vary across datasets. While early positions dominate block selection across datasets, middle and later positions also contribute, indicating that the block selection approaches can retrieve relevant information from throughout the document when necessary.
5.8. Efficiency and Resource Consumption
We compare the reranking time and GPU memory requirements of different models, as shown in Fig. 9. Each experiment is measured on a NVIDIA A100 GPU using the TREC 2019 DL document ranking dataset. As shown in Fig. 9(a), we analyze the reranking time for 100 documents per query (It should be noted the bi-encoder version calculate the block embeddings offline, for fast online search), and in Fig. 9(b), we record the GPU memory usage for various models during training with our settings (see Section 5.3).
(a) Avg. time (seconds) to rerank 100 documets given a query in test set.
(b) GPU memory requirement during training (MB).
Figure 9. Efficiency comparisons on TREC 2019 DL.
5.8.1. Reranking Time
The RankLLaMA model takes the longest time, requiring 14.07 seconds to rerank 100 documents. In contrast, the KeyB2 models were significantly faster. The KeyB2(Llama2)bi model takes 6.25 seconds which is less than half of original time, while the KeyB2(Llama3)cross model require 8.56 seconds. Among all models, the BM25 block matching strategy is the fastest, completing the task in 6 seconds.
5.8.2. GPU Memory Consumption
In terms of GPU memory usage, the RankLLaMA model consumes 33123 MB even with batch size 1, making it the most resource-intensive model. In comparison, The KeyB2 models significantly reduce memory requirements The KeyB2(Llama2)BM25 model used 22523 MB, while the KeyB2(Llama3)cross consume 25297 MB. This represents a substantial memory reduction, ranging from 68% to 76%, demonstrating the efficiency of the KeyB2 models.
These results highlight that KeyB2 models, particularly the KeyB2(Llama2)BM25, offer faster reranking times and lower memory consumption, making them highly efficient and easily deployable for long document retrieval tasks.
5.8.3. Efficiency-Effectiveness Figures
Last, to compare different models’ effectiveness with efficiency together, we draw Fig. 10 to illustrate the overall performance on TREC 2019 DL dataset.
(a) Comparison of speed and effectiveness.
(b) Comparison of memory usage and effectiveness.
Figure 10. Effectiveness with efficiency comparisions on TREC 2019 DL (left and top should be better).
It can be seen the proposed KeyB2 models with different block selection strategies all locate at more left and top parts of the Fig. 10(a) and Fig. 10(b), compared with RankLLaMA model, indicating they have lower reranking latencies, reduced GPU memory requirements, and higher effectiveness than RankLLaMA. If one focuses on the limitation of GPU, one can select to use cross-encoder version. If one focuses on the reranking latency, one can select to use the BM25 or bi-encoder versions, which balance the speed and effectiveness.
@@ 6. Conclusion
In this paper, after exploring attention mechanisms in LLMs for IR and providing insights into the inner workings of decoder-only LLMs, we introduce KeyB2, which incorporates block selection strategies with LLMs to improve long document ranking. By selectively focusing on the most relevant blocks, which is similar to the human relevance judgement process, KeyB2 enhances retrieval accuracy while significantly reducing computational overhead.
Our experiments on various datasets, including TREC 2019 DL document ranking task, Robust04, and MLDR-zh (Chinese), demonstrate that KeyB2 models consistently outperform baseline methods like RankLLaMA, BM25 and KeyB. The models achieve the best improvements with different configurations depending on the dataset: Llama3 with cross-encoder showed the highest gains on English datasets, while the newly proposed bi-encoder configuration delivered superior results on MLDR-zh. Furthermore, our analysis of block selection patterns shows a preference for early document positions across datasets, while demonstrating that KeyB2 effectively retrieves relevant content from various positions as needed, ensuring accurate ranking. Additionally, KeyB2 models show consistent effectiveness in zero-shot scenarios, on the unseen Robust04 dataset. After fine-tuning the performances are better.
Overall, KeyB2 achieves better ranking accuracy while processing fewer tokens than approaches that input entire documents, resulting in faster reranking speed and lower GPU memory usage. These gains make KeyB2 a better solution for high-performance IR task, which is one of the key contributions of this paper: reducing the input size and noise while maintaining the LLMs’ effectiveness is possible by intelligently selecting key blocks from documents à la KeyB. Finally, our design of different strategies successfully improves the performances, pointing to new directions in using LLMs for document retrieval.

### v2

@@ 6.5.2. General Trends
Across all datasets, over half of the top 8 selected blocks come from the first 5 positions in documents, reflecting a common positional bias where initial content is often more relevant to the query. These shows irrelevant noise tokens may be usually after the relevant tokens for these datasets. From Section 3.3, we have shown inserting noise around the relevant document can result in poorer performance. On the contrary, our proposed approach can make the document contexts relevant-focused, with less noise information surrounding, which would improve the attention weights and alignment (e.g., ARAS metric in Section 3), and improve ranking performance. Additionally, the figures indicate that relevance distributions vary across datasets. While early positions dominate block selection across datasets, middle and later positions also contribute, indicating that the block selection approaches can retrieve relevant information from throughout the document when necessary.
@@ 7. Conclusion
We introduced KeyB2, a long-document retrieval framework that first selects salient blocks and then reranks them using decoder-only LLMs. Our attention analysis (RQ1) reveals that when irrelevant tokens are appended, especially near the end of a document, autoregressive LLMs exhibit severe attention dispersion, which undermines relevance estimation. This finding motivates the design of an explicit select-then-combine-then-rerank pipeline with LLMs: KeyB2.
KeyB2 extends the original KeyB paradigm in three ways. First, based on the above analysis and findings from LLM-based IR models, we observe that even LLMs capable of processing longer sequences than BERT still suffer from noise, which hinders attention alignment between relevant parts and the query. This motivates the design of KeyB2. Second, it introduces a next-token-prediction-based scoring mechanism that is compatible with modern decoder-only LLMs. Third, it supports three families of block selectors (BM25, bi-encoder, and cross-encoder), with detailed experimental comparisons, allowing practitioners to trade off between accuracy and efficiency.
Experiments on TREC DL 2019/2023, Robust04, and MLDR-zh demonstrate consistent improvements over RankLLaMA, RankLLaMA-MaxP/AvgP, and earlier KeyB variants. KeyB2 achieves new SOTA results on TREC DL 2019 document reranking (RQ2). The cross-encoder selector achieves the highest scores on English benchmarks, while the bi-encoder variant performs best in MLDR-zh settings. The BM25 selector also performs well on Robust04, which contains documents of moderate length. These results suggest that the choice of selector should be co-tuned with the LLM’s capacity, the selector’s complexity, document length, and deployment constraints (RQ3). Analysis of the selected blocks shows that KeyB2 can surface relevant evidence regardless of its position within the document. This aligns with the findings in Section 3: less noise leads to stronger attention weights and better alignment (e.g., higher ARAS scores) compared to noisy contexts. Moreover, efficiency studies (RQ4) place KeyB2 on the Pareto frontier: it reduces inference time by up to 88% and peak GPU memory usage by over 74% compared to full-document RankLLaMA, while still improving effectiveness.
In short, even in the era of LLMs, selecting key blocks remains a critical component for effective long document ranking. By combining lightweight, explainable selectors with powerful decoder-only LLMs, KeyB2 provides a scalable and adaptable framework that not only enhances ranking performance, efficiency, but also inspires future advances in hybrid long document retrieval paradigms.